Compute the electric potential from a source charge of -2.33 ×10^-13 C if a test charge is 6.45 × 10^-15m away from it


1
Expert's answer
2021-04-29T10:41:05-0400

Solution

Electric potential due to point charge is given

"V=\frac{kq}{r}\\=\frac{9\times10^9\times(-2.33\times10^{-13})}{6.45\times10^{-15})}\\=-3.25\times10^{11}volts"